# Using Half-Precision for GNN Training

### v1

@@ 3. Mixed-Precision GNN Training: Analysis, Related Works, and Overview
Micikevicius et al (micikevicius2017mixed, ) shows that half-precision is as good as a float for diverse kinds of models without any accuracy loss. For this to happen two things need to be ensured during training: 1) as many forward and backward kernel operations can be carried out in half-precision as possible and storing state tensors in half-precision, and 2) the model weight updates must be in float, which automatically happens through mixed-precision module of PyTorch. Currently, no GNN system has applied this successfully. DGL fails to train correctly even on mid-size datasets of Reddit and Ogb-product (Fig. 1(c)), while its half-precision SpMM and SDDMM have not shown any performance improvement. Next, we present root cause analysis for abnormal accuracy and other system design issues, while discussing related works.
3.1. Analysis
@@
We analyze half-precision DGL in more detail and found that a) half-precision SpMM causes value overflow when invoked by mixed-precision training; b) fear of overflow in mixed-precision forces PyTorch kernels to be invoked using float data-type in some cases due to less understanding about system integration to mixed-precision training. A naively designed GNN system in such cases either would invoke float-based sparse kernels or would be performing additional tensor conversion (float to half-precision). DGL does the latter. c) Half-precision kernels do not follow the current state-of-the-practice well-established in dense kernel counterparts. However, we also analyze various float-based GNN systems to show they would face similar issues if extended to half-precision or even borrow the idea from their dense counterparts. We elaborate on these findings next.
@@ 3.1.1. Value Overflow
GCN and GIN use SpMMv whose reduction phase causes output to overflow if a vertex has more neighbors irrespective of the system design. Overflow condition implies that the value range is too small to hold the output in half-precision, resulting in INF (plus or minus infinity). Even modest-sized datasets such as Ogb-Product and Reddit, have many vertices with large enough neighbors. The INF then produces NaN in the follow-up operations, e.g., softmax layer, which involves operation on two INF.
We next show that both the GNN models that rely on SpMMv offer built-in mechanisms to guard against overflow but have not been considered.
GCN. Eq. 1 shows a simplified computation involved in GCN. where, ${\mathbf{D}}$ is the degree matrix of graph ${\mathbf{A}}$. $\mathbf{H}^{(l)}$ is the activation of the previous layer, and $\mathbf{W}^{(l)}$ is the weight matrix of linear layer. The linear layer transforms the activation of the previous layer, which then serves as a dense input matrix to the SpMMv. This also includes a degree-norm scaling. GCN supports 3 different kinds of degree-norm: 1) left: divide the input dense tensor of SpMMv with degree; 2) right: divide the output tensor of SpMMv with degree; and 3) both: divide the input and output dense tensors of SpMMv with the square root of degrees separately which is shown in Eq. 1.
@@
The degree-norm scaling adds stability to feature values. To illustrate, consider the most used degree-norm, right, which simply divides the output features of every vertex produced by SpMMv by its degree. Hence, it is the built-in mechanism that can bring the SpMMv output within the half-precision range. However, the current systems perform degree-norm division after the reduction is over, by then overflow has already happened thus it is unable to protect it.
@@
Some prior float-based GNN works have fused the degree-norm with SpMMv. However, our analysis shows that such approaches are not helpful because they also do degree-norm scaling after the reduction, hence they cannot protect against overflow in half-precision computation. Kindly note that if one uses the left degree-norm, there will be no overflow. However, during backward computation, the degree-norm happens after SpMMv where it is likely to overflow.
@@
GIN. GIN has many variations, we focus on two variations. a) In the most common version, there is no stabilization operation like the degree-norm discussed for GCN. Therefore, there is no method to protect against the value overflow if one uses half-precision. b) DGL offers a few additional reduction operations in addition to the ‘sum’ for the SpMM operation in GIN. One of those options is ‘mean’ which is the same as SpMM followed by degree-norm as discussed in GCN. However, this option exists only in DGL and is implemented the same way as GCN, i.e., the degree-norm is called after SpMM for forward computation. Consequently, this version of GIN is susceptible to the same overflow issue as GCN, while GIN implemented by other float-based GNN systems would always overflow.
3.1.2. Fear of Overflow and Frequent Data-Conversion
This sub-section shows that integration of Custom Kernel System to Mixed-Precision Training has not been clearly understood, thereby half-precision sparse kernels are not invoked as we explain next with a specific example of attention-based GNN, such as GAT, followed by generalizing the observation.
Exponential Operation in GNNs. The final edge-level features in attention-based GNN, also called attention score ($\alpha_{ij}$) are derived using an edge-softmax operator as shown in Eq. 2. Here, the exponential of $e_{ij}$, called attention coefficients, is normalized with the cumulative sum of the exponential of all neighboring $e_{ij}$. $e_{ij}$ gives importance of neighbor $j$’s feature to vertex $i$. These are computed using a series of variants of SDDMM and SpMM kernels. However, PyTorch’s mixed-precision automatically invokes float-based exp operations instead of half-precision computation in Eq. 2.
(2) $\alpha_{ij}=\frac{\exp(e_{ij})}{\sum_{k\in\mathcal{N}_{i}}\exp(e_{ik})}.$
The exponential operation is immediately followed by an SDDMM variant to implement the division operation of Eq. 2 followed by SpMMve. All these follow-up sparse operations receive float type tensors due to this upgrade of just one operation (exponential) to float type. Hence, GNN systems would either a) perform data conversion from float to half-precision to invoke the half-precision sparse kernels as done in DGL or b) be forced to invoke float-based sparse kernels otherwise. The exponential has already performed one data conversion from half to float type of its input tensor.
PyTorch mix-precision stipulates that there is a real chance that half-precision is more likely to overflow when computing individual exponential as it can produce value in the range (0, INF), However, our studies for attention-based GNNs show that exponential operation cannot cause overflow because the model generates a non-positive number for attention coefficient ($e_{ij}$) (computed before Eq. 2). This observation is derived from DGL’s implementation of attention score [^1] https://github.com/dmlc/dgl/blob/master/python/dgl/backend/pytorch/sparse.py##L685, which applies an additive normalization. It first finds the maximum attention coefficient within a neighborhood, then each coefficient subtracts the maximum in that neighborhood, resulting in (-INF, 0] value range. Hence, its exponential is always in the range of (0, 1]. Therefore, the numerator of Eq. 2 can never overflow the half-precision.
This attention score implementation is slightly different than the original GAT model (gat18iclr, ), as the original GAT model has been shown to not converge in some datasets (featgraphsc20, ) even in float precision case.
General Trend. Exponential is not the only operator where mixed-precision automatically upgrades the incoming half-precision tensor to float before computing, but there are a range of operators that are used widely by multiple systems beyond attention-based GNNs that perform this upgrade (PyTorchAMP, ). The list includes common operations such as cross-entropy, log loss, softmax calculation, summation, etc. So almost all GNN should run into these operations, and would not be able to fully utilize half-precision kernels if the system is not designed accordingly. Hence, such system designs cannot be migrated to half-precision computation without performing excessive data-type conversion of tensors or invoking many float-based kernels thereby not providing much benefits in moving to mix-precision training.
3.1.3. Slower Half-Precision Sparse Kernels
We also observe that replacing float-based sparse kernels with half-precision kernels alone is not enough to improve system performance as the currently available half-precision based sparse kernels are slower than their float counterparts as shown earlier in Fig. 1(a) and Fig. 1(b). Specifically, DGL SDDMM, which is open-source, replaces float with the half-precision data-type directly without any system design change.
Based on this observation, we postulate that both the state-of-the-practice and the state-of-the-art for half-precision based GNNs differ from the corresponding dense deep learning model. The state-of-the-practice in half-precision based dense deep learning model focuses on the tensor core, which is well-suited to perform dense matrix multiplication. However, the tensor core is not suited due to the unstructured sparsity present in graph datasets used in GNNs. TC-GNN (wang2023tc, ) which was able to use tensor core in GNN does not perform better than float-based SpMM as shown by prior works (Graphpy, ; GNNBench, ).
Hence, half-precision GNNs must use half2 data-type, whose native support is available in GPUs, for better performance as shown earlier (ho2017exploiting, ) in the dense matrix multiplication before tensor core became prevalent. However, improving the state-of-the-practice by deploying the techniques from dense matrix multiplication (ho2017exploiting, ) alone would not be sufficient to achieve the accuracy and envisioned performance as we explore in this paper.
@@ 3.2. HalfGNN Overview
First of all, HalfGNN adopts the half2-based design and shows the challenges and solutions in adoption. We then show that the resultant design would continue to suffer from the overflow issue identified earlier as the root cause of the value overflow is reduction-related designs of float-based sparse kernels which are not suited directly for half-precision computation. We further show that reduction-related design continues to be the root cause of performance bottlenecks.
@@
To this end, HalfGNN proposes half4 and half8 vector data-type to improve performance for SDDMM by minimizing the inter-thread communication cost which indirectly improves the instruction level parallelism to offer better data-load performance. We further propose discretized reduction scaling to protect against overflow in SpMM by exploiting the built-in degree-norm scaling. The proposed solution is also very well suited to provide workload balancing. We then introduce a staging buffer to remove atomic write, the root cause of performance bottleneck in half-precision SpMM.
@@
The proposed solutions are generic and apply to edge-parallel and vertex-parallel sparse kernels. However, HalfGNN recommends an edge-parallel solution for the best performance.
4. Detailed Design
This section starts with adapting half2 to form a baseline and identifies that using half2 directly in a prior float-based state-of-the-art sparse kernel would not achieve the envisioned accuracy and performance goals without focusing on design around the reduction phase. We then discuss the SDDMM and SpMM design to overcome the issues.
4.1. Adapting Half2 Data-Type
Data-load phase implies bringing the graph and features to GPU registers for computation. This subsection show the challenges and techniques in adopting the half2 data-type.
4.1.1. Better Data-Load using Half2 Data Types
@@
DGL SDDMM does not show any runtime improvement (Fig. 1(b)) as it merely replaces the float data-type with the half data-type. In such an approach, each thread brings one feature (2-byte) from GPU global memory to register for computation, hence a warp performs 64 bytes of data-load in case of full memory coalescing. In comparison, float-based SDDMM issues 128 bytes of data-load. Clearly, to fully utilize the memory transfer capability, it is mandatory to issue more data load to achieve full coalesced memory access. To this end, we adapt the half2 data type, provided natively by CUDA, which has not been used by prior GNN systems.
HalfGNN treats two contiguous half-precision features as one half2 feature, thus allowing a warp to issue 32 half2 values leading to issuing 128 bytes of data-load. Hence, HalfGNN achieves parity with float data-type for data-load performance in terms of size, but two times the number of features. This benefits SDDMM and SpMM both as the features, that we discuss next, are vectors.
@@
We propose a split data-load phase to overcome the challenges associated with adopting half2.
Phase1: Explicitly Loading NZE and Edge-Features. They are explicitly loaded using the edge-parallel method and cached to GPU shared memory so that they can be read again very quickly when needed in the next phase. NZE is of integer data type while edge features, stored in row-major, are fetched sequentially. Hence we can still use vector data type to fetch vector features.
Kindly note that many prior SpMM works (gespmmsc2020, ; huang2021understanding, ; wang2021gnnadvisor, ) allocates a maximum of 32 edges to each warp, hence they cannot achieve better data-load performance for loading edge features, which is in half-precision now. To this end, we propose to allocate at least 64 edges to each warp, hence introducing one more design decision that can introduce better memory coalescing in this phase.
Edge-Feature Mirroring. We notice that only one of the halfs of the edge feature should do dot product with both halfs of half2 vertex features. To illustrate, Fig. 2 (d) for SpMM shows that only $W_{21}$ should be multiplied with {$X_{10}$, $X_{11}$}, but edge-feature data loading brings $W_{21}$ together with $W_{23}$ in our edge-centric design as half2 data-type. Therefore, performing dot product naively would also multiply $W_{23}$ with {$X_{10}$, $X_{11}$}, which is wrong.
Hence, after data loading edge features and before caching them in shared memory, each half is extracted from half2 edge features and is mirrored to create new half2 edge features which are then directly stored in the shared memory for caching. To illustrate, the first half2 word contains {$W_{21}$, $W_{21}$} (to be multiplied with {$X_{10}$, $X_{11}$}), and the next half2 contains {$W_{23}$, $W_{23}$} (to be multiplied with {$X_{30}$, $X_{31}$}). Performing this mirroring after data-load of edge features ensures that edge features can still use the half2 data-type for better data-load performance. Further, computation can be done without any issues or changes.
@@
Phase2: Implicit Vertex Feature Loading. To load vertex-features, threads need to know the vertex ID of non-zero columns, which is already present in the shared memory thanks to phase 1. Vertex-features are stored in rowmajor order, and accessing 2 of them in a group is straightforward. As long as the feature-length is a multiple of 2, a simple type-casting of the features tensor to half2 allows us to use the half2 data type for data-loading. This allows 1 warp thread to carry out data load for 2 features at the same time, instead of 1. So, considering 32 threads in each warp, it can handle 64 features itself. So it doubles the number of features that can be loaded in parallel within each warp.
@@
Feature Padding for Odd Feature Lengths. Kindly note that the feature-length is a configurable parameter in GNN, feature lengths such as 32, 64, etc. are common configurations. However, such lengths could be odd in the last layer of GNNs when the classification category itself is odd. E.g., Reddit has 41 classes to predict. In such cases, we simply use the next even number as class count without any issues.
Figure 3. Reduction in Edge-Parallel SpMM of HalfGNN: Showing 2 CTAs, each containing 4 warps. A warp handles 128 (configurable) edges.
@@
Sub-Warps. We observe that multiple vertex-level features are loaded in a feature-parallel way within a warp. Hence, the threads being used are now half of what would be used in float-based sparse kernels. E.g., when the feature-length is 32, a whole warp processes an edge at any time in the case of float data type. However, using half2, we only need 16 threads of any warp to handle the same features. Hence, we process two non-zero columns at one time by the warp by dynamically creating two sub-warp within a warp. This allows us to use all the threads of the warp leading to better thread utilization. Indeed, we generalize this based on the feature-length, and hence the number of sub-warps within a warp depends on the feature length. E.g., in the case of 16 as feature length, we create 4 sub-warps.
4.1.2. Efficient Computation through Half2 Data-Type
Half arithmetic isn’t natively supported. If one relies on native operators, such as ‘+’ or ‘*’, the data is implicitly converted to float, and the operation is performed in float. The data conversion would be unnecessary. One can use CUDA intrinsic APIs for half computation that avoids this data-type conversion, but it only achieves the same throughput as the float. However, since we already perform data-load in half2, and CUDA intrinsic also provides native support for half2 arithmetic, we use half2 directly for dot product that performs multiplication of two halfs of half2 together. This approach doubles the computation arithmetic throughput.
4.2. SDDMM Reduction: More Vector Data-Types

### v2

@@ 3. Mixed-Precision GNN Training: Analysis, Related Works, and Overview
Micikevicius et al. (micikevicius2017mixed, ) shows that half-precision is as good as a float for training various models without any accuracy loss. They outline two important points during training: 1) many forward and backward kernel operations can be carried out in half-precision, hence state tensors should also be stored in half-precision; 2) the model weight updates must be in float, which automatically happens through the mixed-precision module of PyTorch.
@@
Currently, no GNN system has applied these observations successfully. DGL fails to train correctly even on mid-size datasets of Reddit and Ogb-product (Fig. 1(c)), while its half-precision SpMM and SDDMM have not shown any performance improvement. Next, we present root cause analysis for abnormal accuracy and system performance issues, while discussing related works.
3.1. Analysis and Motivation
@@
This sub-section analyzes half-precision GNN in more detail and shows that a) half-precision sparse kernels significantly differ from corresponding dense deep learning kernels. We also show how the current state-of-the-art sparse kernels are not built on the right hardware abstractions. b) fear of overflow in mixed-precision training forces some PyTorch kernels to be invoked in float data-type due to less understanding about some operations that force additional tensor conversions from float to half and vice versa. c) half-precision SpMM causes value overflow when invoked by mixed-precision training. Naively designed GNN systems in such cases would either invoke float-based sparse kernels or perform additional tensor conversion (float to half-precision). We also analyze various float-based prior GNN systems to show they would face similar issues if extended to half-precision. We elaborate on these findings next.
3.1.1. Slower Half-Precision Sparse Kernels
We postulate that both the state-of-the-practice and the state-of-the-art for half-precision based GNNs differ from the corresponding dense deep learning model and cannot provide performance benefits at this moment.
State-of-the-art half-precision SDDMM by DGL is hardly better than its corresponding float-based SDDMM, as shown earlier in Fig. 1(b). Specifically, it replaces float with the half-precision data type without any system design change, as they follow Fig. 3(a). Hence, it does not lead to any performance gain. On the other hand, DGL uses the Cusparse library for SpMM, and Fig. 1(a) shows that half-precision SpMM is slower than float-based SpMM. Though it is a closed-source code, profiling reveals to a certain extent that it uses atomic instructions to resolve conflicting writes, which is prevalent in many workload-balanced SpMM (section 2.1.3). Atomic write on half-precision is more costly than the corresponding float data type (section 5.2 and  section 6.3.2).
This implies that following a DGL-like strategy of replacing float with half-precision using Fig. 3(a) (or even Fig. 3(b)) approach and/or using atomic instruction to resolve conflicting writes to provide half support in other float-based GNN systems (fu2022tlpgnn, ; huang2021understanding, ; wang2021gnnadvisor, ; wu2021seastar, ; zhang2022understanding, ; gnnone, ; chen2020fusegnn, ; featgraphsc20, ; gespmmsc2020, ; hu2021efficient, ; liang2020engn, ; wang2021flexgraph, ; yan2020hygcn, ; ye2023sparsetir, ; zhang2021boostgcn, ; ferludin2022tf, ) would unlikely result in any performance gain over their float-based design.
State-of-the-practice in half-precision based dense deep learning model focuses on the tensor core, which is well-suited to perform dense matrix multiplication (GeMM). Existing libraries such as CUTLASS by Nvidia are designed to take advantage of tensor cores to perform GeMM. However, the tensor core is unsuited for SpMM due to unstructured sparsity in graph datasets used in GNNs. TC-GNN (wang2023tc, ), which uses tensor cores in GNN, does not perform better than float-based SpMM, as shown by prior works (Graphpy, ; GNNBench, ).
As this paper presents, good practice would be to use half2 data-type in sparse kernels using cuda cores as opposed to tensor cores, whose native support is available in GPUs (Fig. 3c). Earlier, Ho et al. (ho2017exploiting, ) have shown that for the dense matrix multiplication before tensor cores became prevalent. However, this baseline alone, if developed for sparse kernels, would not be sufficient to achieve the accuracy and envisioned performance as discussed in the rest of the paper. Therefore, this paper complements approaches such as CUTLASS libraries as we focus on sparse computations.
3.1.2. Invoking Half-Precision Sparse Kernels with Frequent Data-Conversion
This sub-section shows that invocation of half-precision sparse kernel in mixed-precision training environment has not been clearly understood, thereby half-precision sparse kernels may not be invoked or may be invoked with an additional data-conversion as we explain next with a specific example of attention-based GNN, such as GAT, followed by generalizing the observation.
Exponential Operation in GNNs. The final edge-level features in attention-based GNN, also called attention score ($\alpha_{ij}$), are derived using an edge-softmax operator. It represents the relative importance of neighbor $j$’s feature to vertex $i$. For numerical stability, it is computed using Eq. 1. Here $e_{ij}$ is the initial feature on each edge, which itself is derived from the features of its source and destination vertices using trainable parameters and an SDDMM variant. It then uses a SpMM variant to get each neighborhood’s maximum edge feature ($m_{i}$).
(1) $m_{i}={\max_{j\in\mathcal{N}_{i}}(e_{ij})},e^{\prime}_{ij}=\exp(e_{ij}-m_{i}),\alpha_{ij}=\frac{e^{\prime}_{ij}}{\sum_{j\in\mathcal{N}_{i}}(e^{\prime}_{ij})}$
However, PyTorch’s mixed-precision automatically invokes float-based exp in Eq. 1 operations instead of half-precision computation. The exponential operation is immediately followed by an SDDMM variant to implement the division operation of Eq. 1. The output is then sent to SpMMve (not shown in the Eq.) Hence, all these follow-up sparse operations receive float tensors due to this upgrade of just one operation (exponential) to float type. Hence, GNN systems would either a) perform data conversion from float to half-precision to invoke the half-precision sparse kernels as done in DGL or b) be forced to invoke float-based sparse kernels otherwise. The exp has already performed a data conversion from half to float on its input.
Fear of Overflow. PyTorch mix-precision stipulates that there is a real chance that half-precision is more likely to overflow when computing individual exponential as it can produce values in the range (0, INF). However, we point that $e_{ij}-m_{i}$ can only be a non-positive number in the range of (-INF, 0], hence its exponential cannot cause overflow as the result would only be in (0, 1] which prior works do not take advantage of.
General Trend. Exponential is not the only operator where mixed-precision automatically upgrades the incoming half-precision tensor to float before computing. Many operators used widely by GNN systems perform this upgrade (PyTorchAMP, ). The list includes common operations such as cross-entropy, log loss, softmax calculation, summation, etc. So, almost all GNNs should run into these operations where invoking a half-precision sparse kernel would entail a data conversion if the system is not designed accordingly. Even for GCN, CSR-based design of pytorch-geometric (fey2019fast, ) shows the data conversion problem where many operations result in half-precision to float data conversion, and subsequent operations are done in float. Its earlier COO-based design often runs out of memory in Reddit, OGB-product, and other mid-size datasets (Graphpy, ; chen2020fusegnn, ; huang2021understanding, ), hence, it has not been analyzed.
@@ 3.1.3. Value Overflow
GCN and GIN use SpMMv whose reduction phase causes output to overflow if a vertex has more neighbors, irrespective of the system design in half-precision. Overflow condition implies that the value range is too small to hold the output in half-precision, resulting in INF (plus or minus infinity). Even modest-sized datasets, such as Ogb-Product and Reddit, have many vertices with large enough neighbors. The INF then produces NaN in the follow-up operations, e.g., the softmax layer, which operates on two INF. Hence, approaches of Fig. 3(b) and (c) would always result in value overflow in SpMM when neighbors are large. However, if one chooses to use Fig. 3(a), SpMM would still overflow during reduction as it would likely produce a float value that overflows when converted to half-precision.
Next, we show that GNN models offer built-in mechanisms to guard against overflow but have not been used by prior works.
GCN. Eq. 2 shows its simplified computation. where, $\mathbf{D}$} is the degree matrix of graph $\mathbf{A}$}. $\mathbf{H}^{(l)}$ is the activation of the previous layer, and $\mathbf{W}^{(l)}$ is the weight matrix of linear layer. The linear layer transforms the activation of the previous layer, which then serves as a dense input matrix to the SpMMv. This also includes a degree-norm scaling. GCN supports 3 different kinds of degree-norm: 1) left: divide the input dense tensor of SpMMv with degree; 2) right: divide the output tensor of SpMMv with degree; and 3) both: divide the input and output dense tensors of SpMMv with the square root of degrees separately, which is shown in Eq. 2.
@@
The degree-norm scaling adds stability to feature values. To illustrate, consider the frequently used degree-norm, right, which divides the output vertex features of SpMM by the vertex degree. Hence, the degree-norm can bring the SpMMv output within the half-precision range. However, the current systems perform degree-norm after the reduction is over, by then overflow has already happened. Some prior float-based GNNs (wang2021gnnadvisor, ) have fused the degree-norm with SpMMv. However, our analysis shows that such approaches are not helpful because they also do degree-norm after the reduction phase. Hence, prior works cannot protect against overflow in half-precision computation.
@@
Kindly note that if one uses the left degree-norm, there will be no overflow. However, during backward computation, the degree-norm happens after SpMMv (right), where it is likely to overflow.
@@
GIN. GIN has many variations, we focus on two variations. a) In the most common version, there is no stabilization operation like the degree-norm discussed for GCN. Therefore, there is no method to protect against the value overflow if one uses half-precision. b) DGL offers a few additional reduction operations in addition to the ‘sum’ for the SpMM operation in GIN. One of those options is ‘mean’ reduction, which is the same as SpMM followed by degree-norm as discussed in GCN. However, this DGL option is implemented the same way as GCN, i.e., the degree-norm is called after SpMM for forward computation. Consequently, this version of GIN is susceptible to the same overflow issue as GCN, while GIN implemented by other float-based GNN systems would always overflow.
@@ 3.2. HalfGNN Overview
The proposed HalfGNN tackles the performance and accuracy issues together. First, HalfGNN adopts the half2-based design to arrive at a better half-precision baseline and shows unique challenges in its adoption. It proposes edge-feature mirroring, feature padding for odd vertex features, and introducing sub-warps to impart better thread utilization to adopt half2. However, reduction-related design, including reliance on atomic write, continues to cause performance bottlenecks. Lastly, the resultant design would still suffer from the overflow issue, whose root cause is reduction-related designs.
@@
To this end, HalfGNN further proposes half4 (64 bits) and half8 (128 bits) vector data types, packing four and eight half-precision data respectively. These two vector types natively support data loading. Utilizing them improves SDDMM performance by minimizing the inter-thread communication cost, which indirectly improves the instruction-level parallelism to offer better data-load performance. For SpMM, discretized reduction scaling is proposed where only a finite batch of neighbors is reduced and degree scaled before moving to the next batch of neighbors. The design allows us to implement multiple workload-balancing solutions. We then introduce a staging buffer to remove atomic write, the root cause of the performance bottleneck in half-precision SpMM. This design also protects against overflow in SpMM due to the discrete nature of reduction and degree-norm scaling.
@@
The proposed solutions are generic and apply to both the edge-parallel and the vertex-parallel sparse kernels. E.g., the discretized reduction is very well suited to provide workload balancing in edge-centric and vertex-centric sparse kernels. However, HalfGNN recommends an edge-parallel solution for the best performance, which we discuss in detail, while we also show its applicability in a vertex-parallel solution(section 6.3) of Huang et al. (huang2021understanding, ).
4. Adapting Half2 Data-Type
This section focuses on adapting half2 to form a baseline, which we can analyze to improve performance and accuracy. We discuss the challenges and techniques in adopting the half2 data type. We strive to incorporate state-of-the-art techniques from float-based sparse kernel design, such as workload balancing, while developing the baseline. However, the focus is also to show the applicability of proposed techniques to edge-parallel and vertex-parallel design paradigms to claim general applicability.
4.1. Better Data-Load using Half2 Data Types
@@
Substituting half in place of float would always compromise the data load performance. In such an approach, each thread brings one feature (2-byte) from GPU global memory to register for computation, hence a warp performs only 64 bytes of data-load. In comparison, a float-based system issues 128 bytes of data load per warp, which achieves full coalesced memory access. This is evident from the DGL SDDMM result (Fig. 1(b)) that does not show any runtime improvement as it merely replaces the float with the half data type.
@@
To fully utilize the memory transfer capability, HalfGNN adapts the half2 data type, provided natively by CUDA, which prior GNN systems have not used. We propose a split data-load phase to overcome the challenges associated with adopting half2.
4.1.1. Phase1: Explicitly Loading NZE and Edge-Features.
We observe that each NZE and its edge-feature are used $|F|$ times (feature length of the vertices) during computation. To illustrate, the role of NZE is to know the column (vertex ID) so that its vertex features (vector) could be fetched for SpMM and SDDMM. As different warp threads fetch different features of the same column, the column ID could be used by them if we fetch it once and cache it. Similarly, the edge-feature of the NZE performs a dot product with all the corresponding vertex features in SpMM.
However, in a naive feature-parallel method, each thread of the warp would load the same NZE and edge-feature, leading to repeated and non-coalesced access. HalfGNN deploys an explicit edge-parallel method to load and cache both of them to the GPU shared-memory. This allows their data-load to be performed in a coalesced fashion, while caching in shared memory achieves the effect of broadcasting, allowing feature-parallel threads to read locally when needed in the next phase.
Further, HalfGNN treats two contiguous half-precision edge features as one half2 feature. A simple type-casting is all that is needed to cast the half features to the half2 data type, if designed carefully (see feature padding for challenges involved). Thus, a warp can issue 32 half2 values, leading to 128 bytes of data-load. To take advantage of this phase, at least 64 edges must be allocated to each warp to process. This would be equivalent to 32 half2 (128 bytes) features.
4.1.2. Phase2: Implicit Vertex Feature Loading.
@@
To load vertex features, threads need to know the column ID of each NZE, which is already present in the shared memory thanks to phase 1. Vertex features are stored in row-major order, and accessing 2 of them in a group is straightforward. As long as the feature length is a multiple of 2, a simple type-casting of the features tensor to half2 allows us to use the half2 data type for data-loading. This allows 1 warp thread to carry out data load for 2 features at the same time, instead of 1. So, considering 32 threads in each warp, it can handle 64 features. So, it doubles the number of features that can be loaded in parallel within each warp over a float-based system.
@@
Feature Padding for Odd Feature Lengths. Kindly note that the feature length is a configurable parameter in GNN, feature lengths such as 32, 64, etc., are common configurations. However, such lengths could be odd in the last layer of GNNs when the classification category itself is odd. E.g., Reddit has 41 classes to predict. In such cases, a simple type-casting half to half2 does not work due to memory alignment issues, where hardware would not allow accessing half2 values whose address is not a multiple of 4 bytes. In such a case, we simply use the next even number as the class count without any issues.
@@
Sub-Warps. We observe that vertex-level features are loaded in a feature-parallel way within a warp. Hence, the threads being used are now half of what would be used in float-based sparse kernels. E.g., when the feature-length is 32, we only need 16 threads of any warp to when using half2,. Hence, we process two non-zero columns at one time by the warp by dynamically creating two sub-warps within a warp. This allows us to use all the threads of the warp, leading to better thread utilization. Indeed, we generalize this based on the feature-length, and hence the number of sub-warps within a warp depends on the feature length. E.g., in the case of 16 as feature length, we create 4 sub-warps.
4.2. Faster Computation using Half2 Data-Type
Fig. 3(a) approach illustrates that native arithmetic using operators such as ‘+’ or ‘*’ causes implicit data conversion of half-precision operands to float, and the operation is performed in float. Similarly, CUDA intrinsic APIs for half computation avoid this data-type conversion but achieve the same throughput as the float (Fig. 3(b)). Our approach is to follow Fig. 3(c) as CUDA intrinsic natively supports half2 arithmetic, while our data is already in half2. This approach doubles the arithmetic computation throughput but faces a unique challenge, discussed next.
Edge-Feature Mirroring. The dot-product phase in SpMM multiplies the edge feature (in half) by the vertex features (in half) of the non-zero column of the edge. However, usage of half2 packs two edge features as one half2. This means that only one of the halfs of the edge feature should do the dot product with both halfs of the half2 vertex features for any given edge. To illustrate, Fig. 2 (d) for SpMM shows that only $W_{21}$ should be multiplied with $X_{10}$, $X_{11}$}, but edge-feature data loading brings $W_{21}$ together with $W_{23}$ in our edge-centric design as half2 data-type. So, performing dot product naively would also multiply $W_{23}$ with $X_{10}$, $X_{11}$}, which is wrong.
Hence, in phase-1 of data-load (section 4.2), after data loading edge features and before caching them in shared memory, we extract each half from half2 edge features, and is mirrored to create new half2 edge features which are then directly stored in the shared memory for caching. To illustrate, the first half2 word contains $W_{21}$, $W_{21}$} (to be multiplied with $X_{10}$, $X_{11}$}), and the next half2 contains $W_{23}$, $W_{23}$} (to be multiplied with $X_{30}$, $X_{31}$}).
Performing this mirroring does not introduce any additional data-load as it happens after the data-load of edge features. Hence, it ensures that edge features can still use the half2 data type for the data-load and dot product phases. The mirroring allows us to utilize the approach of Fig. 3 (c) resulting in doubling the arithmetic throughput to perform the dot product compared to float-based computation.
Figure 4. Reduction in Edge-Parallel SpMM of HalfGNN showing 2 CTAs, each containing 4 warps. Each warp handles 128 edges. Vertex feature length is 64 ($|F|$). a) The dot products are locally aggregated by each warp by performing a running reduction. b) The non-conflicting reduced output is written directly in the output tensor ($Y$). c) The conflicting reduced output ($|F|$ per warp) is first aggregated within CTA using shared memory, which generates many non-conflicting writes and one final conflicting write ($|F|$ per CTA). The latter is sent to the staging buffer, which is handled by a follow-up kernel to write them to the output tensor ($Y$). See section 5.2.3 for details on the method.
5. Reduction Centric Design
The developed baseline using the half2 data type still does not achieve the envisioned performance and accuracy goals without focusing on a design centered around the reduction phase. This section discusses the SDDMM and SpMM design to overcome the issue, which changes the data load and computation phases to accommodate the concerns arising from their reduction design.
5.1. SDDMM: More Vector Data-Types
SDDMM reduction happens on the dot products along the feature dimension. As feature parallelism is used to load features and compute dot products, the reduction requires inter-thread communication to generate a half feature per NZE (edge).
5.1.1. Observation